# arXiv:2603.01914v1[cs.CL] 02 Mar 2026

### v1

@@ Abstract
Test-time scaling via recurrent/iterative Transformers enables large language models to spend more computation at inference, but most pretrained recurrent LMs run a fixed number of iterations, wasting compute on easy tokens and lacking token-wise adaptivity. Following the core idea of Adaptive Computation Time(ACT) and Early Exit(EE), we propose AdaPonderLM, a self-supervised recurrent language model that learns token-wise early exiting during pretraining without manually tuned per-token/per-layer pruning ratios. AdaPonderLM uses iteration-specific MLP gates with a monotonic halting mask to decide when each token stops recurring, and introduces a KV reuse mechanism that reuses cached key/value states for halted tokens, ensuring train–test consistency and practical acceleration. Across Pythia backbones from 70M to 410M (pretraining) and up to 2.8B (continued pretraining), AdaPonderLM reduces inference compute at about 10% while maintaining comparable language modeling perplexity and competitive downstream accuracy. Our analysis shows the learned gates allocate more computation to high-NLL (hard) tokens, exhibiting adaptive computation time behavior in a fully self-supervised setting. Meanwhile, under iso-FLOPs, the learned halting policy consistently outperforms fixed pruning, showing AdaPonderLM allocates compute to the right tokens rather than just reducing average depth.
@@ 1 Introduction
Following the core idea of Adaptive Computation Time (ACT) [graves2016adaptive], we highlight Early Exit (EE) as a practical way to alleviate the inefficiency of fixed-depth recurrence. EE has been widely studied in CNNs and Transformers to reduce computation by skipping layers [teerapittayanon2016branchynet, bolukbasi2017adaptive, xin2020deebert, schwartz2020right]. In decoder-only architectures, methods such as Universal Transformer [dehghani2018universal] and LayerSkip [elhoushi2024layerskip] further demonstrate the potential of early exiting. In recurrent architectures, however, the same principle is particularly natural for enabling temporal adaptivity: rather than traversing a predetermined number of recurrent steps, the model can learn to halt its recurrence dynamically. Concretely, specialized halting gates or confidence-based criteria can determine an input-dependent termination time, thereby decoupling computational cost from parameter count and allocating variable amounts of “thinking time” to tokens according to their inherent difficulty [graves2016adaptive, bengio2015conditional].
@@ 3.1 MLP-based Gate: Model Framework
Let the embedding matrix for vocabulary be $V=[e_{1},e_{2},\ldots,e_{|\mathcal{V}|}]\in\mathbb{R}^{|\mathcal{V}|\times d}$. Given an input token sequence of length $n$, we obtain the initial embeddings $E^{0}=[e^{0}_{1},e^{0}_{2},\ldots,e^{0}_{n}]$, where $e^{0}_{t}\in\mathbb{R}^{d}$. We then run a recurrent refinement process indexed by iteration superscripts $i$.

### v2

# AdaPonderLM: Gated Pondering Language Models with Token-Wise Adaptive Depth

@@ Abstract
Test-time scaling via recurrent/iterative Transformers enables large language models to spend more computation at inference, but most pretrained recurrent LMs run a fixed number of iterations, wasting compute on easy tokens and lacking token-wise adaptivity. Following the core idea of Adaptive Computation Time(ACT) and Early Exit($E$), we propose AdaPonderLM, a self-supervised recurrent language model that learns token-wise early exiting during pretraining without manually tuned per-token/per-layer pruning ratios. AdaPonderLM uses iteration-specific MLP gates with a monotonic halting mask to decide when each token stops recurring, and introduces a KV reuse mechanism that reuses cached key/value states for halted tokens, ensuring train–test consistency and practical acceleration. Across Pythia backbones from 70M to 410M (pretraining) and up to 2.8B (continued pretraining), AdaPonderLM reduces inference compute at about 10% while maintaining comparable language modeling perplexity and competitive downstream accuracy. Our analysis shows the learned gates allocate more computation to high-NLL (hard) tokens, exhibiting adaptive computation time behavior in a fully self-supervised setting. Meanwhile, under iso-FLOPs, the learned halting policy consistently outperforms fixed pruning, showing AdaPonderLM allocates compute to the right tokens rather than just reducing average depth.
@@ 1 Introduction
Following the core idea of Adaptive Computation Time (ACT) [graves2016adaptive], we highlight Early Exit ($E$) as a practical way to alleviate the inefficiency of fixed-depth recurrence. $E$ has been widely studied in CNNs and Transformers to reduce computation by skipping layers [teerapittayanon2016branchynet, bolukbasi2017adaptive, xin2020deebert, schwartz2020right]. In decoder-only architectures, methods such as Universal Transformer [dehghani2018universal] and LayerSkip [elhoushi2024layerskip] further demonstrate the potential of early exiting. In recurrent architectures, however, the same principle is particularly natural for enabling temporal adaptivity: rather than traversing a predetermined number of recurrent steps, the model can learn to halt its recurrence dynamically. Concretely, specialized halting gates or confidence-based criteria can determine an input-dependent termination time, thereby decoupling computational cost from parameter count and allocating variable amounts of “thinking time” to tokens according to their inherent difficulty [graves2016adaptive, bengio2015conditional].
@@ 3.1 MLP-based Gate: Model Framework
Let the embedding matrix for vocabulary be $V=[e_{1},e_{2},\ldots,e_{|\mathcal{V}|}]\in\mathbb{R}^{|\mathcal{V}|\times d}$. Given an input token sequence of length $n$, we obtain the initial embeddings $E^{0}=[e^{0}_{1},e^{0}_{2},\ldots,e^{0}_{n}]$, where $e^{0}_{t}\in\mathbb{R}^{d}$. We then run a recurrent refinement process indexed by iteration superscripts ii.